# Bidding in Ancillary Service Markets: An Analytical Approach Using Extreme Value Theory

### v1

@@ I-A Background and Motivation
Flexibility aggregators can pool and manage distributed energy resources, providing services such as frequency regulation and load balancing to support grid stability. This study focuses on electric vehicle (EV) aggregators that participate in frequency-supporting ancillary service markets in the Nordic region [1]. Due to the stochastic nature of EV charging behavior and availability, ensuring reliable participation in these markets presents significant challenges, requiring strategies to manage uncertainty in the aggregator’s flexibility supply.
To facilitate the participation of flexible yet stochastic resources in ancillary service markets, the Danish transmission system operator (TSO), Energinet, has introduced grid codes that establish a minimum 90% reliability requirement for the full availability of ancillary service bids [2]. By explicitly setting a failure threshold of 10%, which allows for partial or full unavailability of the ancillary service bid, this requirement encourages resources with uncertain consumption or production baselines to bid their flexibility. This, in turn, increases the supply of such services, potentially reduces service prices, and lowers the TSO’s procurement costs. However, accommodating this uncertainty in service availability during activation events may compromise grid security.
This paper takes the aggregator’s perspective to explore how to exploit this requirement. The investigation of its grid security implications, as well as the extent to which the TSO must purchase additional volumes of services to ensure grid security at a desired confidence level, is left for future work. To meet the 90% reliability requirement, aggregators must employ optimization techniques that incorporate probabilistic constraints to manage uncertainty while ensuring compliance with TSO requirements. Chance-constrained optimization is a natural choice for this task.
The paper addresses the bidding strategy of an EV aggregator using the minimum 90% reliability requirement to place hourly reserve capacity bids in the Nordic ancillary service markets. Without loss of generality, we focus on the market for frequency containment reserve for disturbances (FCR-D). Bidding decisions are made in the day-ahead stage, with uncertain future consumption from the EV aggregation. The main challenge is modeling uncertainties in EV charging patterns while adhering to the TSO requirement. We also explore how the aggregator’s strategy would change if the TSO adjusts the requirement. As our main methodological contribution, we develop a chance-constrained optimization model using concepts from extreme value theory and propose an analytical reformulation. This contrasts with the commonly used sample-based approach, which would require an increasing number of samples to model the underlying uncertainty.
@@ I-B Literature Review
A common approach to solving optimization problems with uncertainty is scenario-based methods, where uncertainties are represented by a finite set of scenarios sampled from the underlying probability distribution or empirical data. While straightforward, these methods often require a large number of scenarios to accurately capture the underlying uncertainty, leading to increased computational costs [6, 7]. An alternative approach involves approximation techniques or analytical reformulations for specific types of probability distributions. These methods convert optimization problems with probabilistic constraints into analytically tractable forms, enabling more efficient solution methods [11, 8, 9, 10, 12].
@@ I-C Contributions
This study develops a chance-constrained optimization model for an EV aggregator in Nordic ancillary service markets, accounting for the stochastic consumption baseline of EV fleets. By applying concepts from extreme value theory to analyze the tail behavior of EV charging patterns, the model focuses on rare, high-impact events that could lead to constraint violations and, consequently, unsuccessful flexibility service delivery. This approach improves risk characterization and ensures the model effectively handles worst-case scenarios. An analytical solution is provided that is less computationally demanding compared to commonly used sample-based methods. Additionally, the study explores how varying failure thresholds set by Energinet impact the ancillary service market participation of the EV aggregator, offering practical insights for both aggregators and the TSO on how failure thresholds influence bidding strategies and ancillary service supply.
@@ II Preliminaries
Figure 1: Illustrative examples of flexibility estimation: The figure on the left illustrates how downwards and upwards flexibility for the aggregated case are calculated. Downwards flexibility ($r^{\downarrow}$) is determined as the difference between the lowest point of the maximum consumption of the EVs and the highest actual consumption across the hour, while upwards flexibility ($r^{\uparrow}$) is the minimum of the actual consumption values during the hour. The middle and right figures show how energy flexibility ($r^{\rm{\rm{E_{20}}}}$) is calculated for a single EV and for the aggregated case. The single EV case represents a complete charging session, while the aggregated case captures a time interval in the day with large fluctuations in available flexibility, typically during the morning when many charging sessions end. Energy flexibility is calculated based on the current state of charge (SoC) and the amount of power that can be applied to the EV over the next 20 minutes without exceeding the battery’s capacity. For instance, at hour 1 in the single EV case, the SoC is 30 kWh, and the battery capacity is 75 kWh, meaning that, theoretically, the charge box (CB) could apply a maximum of 45 kWh in the next 20 minutes to fully charge the EV. However, due to the CB’s physical constraint of 12 kW, the actual available energy flexibility is much lower. Similarly, in the aggregated case, the theoretical and realistic energy flexibilities are on vastly different scales, reflecting the system’s capacity and constraints.
@@ II-B The P90 Requirement
To maintain grid stability, the provision of ancillary services must be highly reliable. However, requiring a 100% reliability for ancillary service bids would restrict participation from many non-conventional technologies, such as wind power or aggregated flexible loads like EVs. To enable stochastic flexible resources to participate in the FCR markets, including FCR-D which is the focus of this paper, Energinet has introduced the P90 requirement, as outlined in [2] and [5].
The P90 requirement imposes that flexibility providers, such as EV aggregators, can participate in the FCR markets, provided that their reserved capacity bid is fully realized at least 90% of the time. A flexibility provider, after being qualified by the TSO, can participate in the relevant ancillary service markets, from which point Energinet will continuously check ex-post whether a reservation capacity bid by the participant truly exists or not. This is done no matter if the reserve is activated, as an extra measure of security of supply. For example, an upwards reserve bid of 10 kW will be seen as unavailable if the true consumption realization of an EV aggregator is 8 kW only. An unavailable bid, partially or fully, counts as a reserve shortfall. Energinet conducts these checks at appropriate time intervals, e.g., every quarter, and in the case of violation of the P90 requirement, the provider will be excluded from further participation in the market and must re-apply for the qualification. This requirement can be expressed mathematically using chance constraints of the form $\mathbb{P}(\cdot)\geq 1-\varepsilon$, where $\varepsilon$ represents the maximum 10% allowance for failure in successfully delivering the service.
@@ II-C The LER Requirement
In addition to the P90 requirement, Energinet has imposed an additional requirement for resources with limited energy reservoirs (LER), such as batteries and EV aggregators. The primary focus of LER requirements is to ensure that resources with limited energy capacity can maintain energy availability despite their limitations. This includes adhering to energy management protocols and ensuring the ability to provide continuous activation for a certain minimum time duration when needed. The definition of this requirement is outlined in [2] and [5].
The specifics of the LER requirement vary between the FCR-N and FCR-D services. Focusing on FCR-D services in this paper, the requirement imposes that for an LER unit, when bidding in one direction of the FCR-D market (either up or down), at least 20% of the reserve bid must also be available in the opposite direction. For EV aggregators, this LER requirement applies only in the downwards direction, as the activation (i.e., increasing power consumption) is constrained by the battery energy storage capacity of EVs. In contrast, activation in the upwards direction (i.e., reducing power consumption) is not constrained by the battery capacity and is treated similarly to other non-LER flexible loads. Additionally, the requirement mandates that the service should be available for at least 20 continuous minutes once activated. These conditions result in the following constraints on the upwards and downwards reserve capacity bids of the EV aggregator (in kW), denoted as $b^{\uparrow}$ and $b^{\downarrow}$, to be submitted to the FCR-D market:
@@
where $\{R^{\uparrow},R^{\downarrow},R^{\rm{E_{20}}}\}$ are the three random variables representing the available upwards, downwards, and energy flexibility, respectively. Downwards flexibility, $R^{\downarrow}$, refers to the potential for increasing consumption of the grid-connected EV fleet, while upwards flexibility $R^{\uparrow}$, refers to the potential for reducing consumption. In addition, energy flexibility, $R^{\rm{E_{20}}}$, is calculated based on the current state of charge of the entire EV fleet and the amount of power that can be applied to the fleet over the next 20 minutes without exceeding their battery capacity. The realizations of these random variables are denoted by $\{r^{\uparrow},r^{\downarrow},r^{\rm{E_{20}}}\}$. A thorough explanation of how these realizations are defined can be found in Appendix A, along with a description of the case study. For readers seeking a quicker overview, Fig. 1 provides graphical representations of how the flexibilities are estimated. Specifically, the realizations $\{r^{\uparrow},r^{\downarrow},r^{\rm{E_{20}}}\}$ are defined as the extreme values of flexibility within the hour, which in our case corresponds to the minimum available flexibility during the hour. This approach contrasts with [5], where similar sets are defined using minute-level realizations. By taking the minimum value over the entire hour, we reduce the computational burden, as the number of data points is decreased by a factor of 60.
@@ III Chance-Constrained Bidding Model
To incorporate the requirements imposed by Energinet into the bidding strategy, we use chance-constrained optimization. This approach handles uncertainty by ensuring that certain constraints are satisfied with a specified probability, rather than with absolute certainty. For EV aggregators participating in ancillary service markets, chance constraints allow them to make bids that balance reliability and resource utilization. Specifically, Energinet’s grid code requires that aggregators commit to a service provision that is fully available at least 90% of the time, allowing for a 10% margin of non-compliance due to uncertainties in EV availability.
Following the P90 and LER requirements, we adopt a joint chance-constrained optimization model similar to the one proposed in [5]:
@@
where the objective function (2a) maximizes the combined reserve capacity bids into the FCR-D up and down markets. Additionally, (2b) defines the joint chance constraint that incorporates both the LER and P90 requirements. Specifically, the chance constraint $\mathbb{P}(\cdot)\geq 1-\varepsilon$ ensures that the probability of jointly satisfying the constraints within the corresponding hour is at least $1-\varepsilon$. According to the P90 requirement, $\varepsilon=0.1$.
@@
Chance-constrained programs are generally computationally intractable. However, if the underlying probability distribution has certain properties, analytical reformulations of the chance constraints can be developed. For example, analytical methods like the Bonferroni approximation [14] transform probabilistic constraints into a series of more tractable constraints, simplifying the problem at the potential cost of precision. Alternatively, sample-based methods represent uncertainty through a set of likely scenarios, allowing for optimization that accounts for a range of possible outcomes. These approaches enable EV aggregators to optimize their bidding strategies while managing the stochastic nature of their resources within the operational constraints of ancillary service markets.
@@
In Fig. 2 scatter plots of the in-sample and out-of-sample data used in our case study are shown. In-sample data refers to the set used in the optimization problem to determine optimal bids, while out-of-sample data is used ex-post for bid evaluation. Extreme value theory [15] focuses on the behavior of extreme values (maxima or minima) within a dataset, particularly as the sample size increases. We are especially interested in the extreme values, which in our case correspond to those below the dotted line in Fig. 2, representing the empirical 10th percentile of the in-sample data. To avoid bias, we solve the model in multiple runs, each with different sets of in-sample and out-of-sample data, while keeping the number of samples in each set unchanged.
@@
To derive an analytical solution to (2), we aim to fit a distribution to the tail data (samples below the empirical 10th percentile) that allows for an analytical reformulation of the chance constraint. Among the various options for modeling tail data, we tested both the Weibull and Generalized Pareto distributions. We observed comparatively better out-of-sample performance when fitting the Weibull distribution.
@@
It is worth noting that in the literature on extreme value theory, the Generalized Extreme Value (GEV) distribution is more commonly used as it emerges as the limiting distribution of the maximum of $n$ independent and identically distributed random variables when $n$ grows large. However, the convergence to this limit law may be poor. In addition, the GEV distribution introduces additional parameters that require tuning, which also may require more samples.
@@ IV-A Fitting a Weibull Distribution to the Tail Data
Aiming to obtain an analytical solution, we fit a Weibull distribution to the tails of our data. The rationale for choosing this distribution lies in its flexible shape parameters, allowing to capture a wide range of tail behaviors. This flexibility is particularly advantageous, as this distribution can effectively model both skewed and heavy-tailed data. Its ability to adapt to both light and heavy tails enables us to accurately capture the realistic variability and potential extremes in the data.
The Weibull distribution has extensive applications in fields such as reliability engineering, where it models failure times and lifetimes, and in risk management for assessing minimum extreme events.
@@
which we maximize over a range of $\gamma$-values. Looking back at Fig. 3, the Weibull distribution is fitted to the lower tail of the sampled data, with the shape parameter $\hat{\gamma}=0.9265$.
As we will require later, we also define the cumulative distribution function (CDF) of the Weibull distribution as:

### v2

@@ I-A Background and motivation
Flexibility aggregators can pool and manage smaller flexible units located on the consumer side of power systems, known as distributed energy resources. These resources can provide services such as frequency regulation and load balancing, contributing to the stability of the electrical grid. This study focuses on electric vehicle (EV) aggregators that participate in frequency-supporting ancillary service markets in the Nordic region [1]. However, the stochastic nature of the charging behavior and availability of electric vehicles creates significant challenges in ensuring reliable market participation, which requires strategies to manage uncertainty in the flexible supply of the aggregator.
To enable the participation of flexible yet stochastic resources in ancillary service markets, the Danish transmission system operator (TSO), Energinet, has introduced grid codes requiring a minimum 90% reliability for the full availability of ancillary service bids [2]. By explicitly allowing up to 10% failure, meaning partial or complete unavailability of the bid, this requirement encourages participation from resources with uncertain consumption or production baselines. As a result, it increases the supply of ancillary services, potentially reduces service prices, and reduces the TSO’s procurement costs. However, accommodating uncertainty in service availability during activation events may compromise grid security.
This paper examines how aggregators can leverage this requirement. The analysis of its implications for grid security and the additional service volumes the TSO may need to procure to maintain a desired confidence level is left for future work. To meet the 90% reliability requirement, aggregators must apply optimization techniques that incorporate probabilistic constraints to manage uncertainty while ensuring compliance with the TSO regulations. Chance-constrained optimization naturally lends itself to this task.
We focus on the bidding strategy of an EV aggregator that uses the 90% reliability requirement to place hourly reserve capacity bids in the Nordic ancillary service markets. Without loss of generality, we consider the market for frequency containment reserve for disturbances (FCR-D). Bidding decisions are made on the day-ahead stage, where future EV consumption remains uncertain. The key challenge is modeling these uncertainties while ensuring compliance with the TSO requirement. Furthermore, we explore how the aggregators strategy would change if the TSO adjusts the reliability threshold.
As our main methodological contribution, we develop a chance-constrained optimization model using concepts from extreme value theory and propose an analytical reformulation. This contrasts with the widely used sample-based approach, which requires an increasing number of samples to capture the underlying uncertainty.
@@ I-B Literature review
A common approach to solving optimization problems with uncertainty is scenario-based methods, where uncertainties are represented by a finite set of scenarios sampled from the underlying probability distribution or empirical data [6]. While straightforward, these methods often require a large number of scenarios to accurately capture the underlying uncertainty, leading to increased computational costs [7, 8]. An alternative approach involves approximation techniques or analytical reformulations for specific types of probability distributions. These methods convert optimization problems with probabilistic constraints into analytically tractable forms, enabling more efficient solution methods [9, 10, 11, 12, 13].
@@ I-C Contributions
This study develops a chance-constrained optimization model for an EV aggregator in Nordic ancillary service markets, accounting for the stochastic consumption baseline of EV fleets. By applying concepts from extreme value theory to analyze the tail behavior of EV charging patterns, the model focuses on rare, high-impact events that could lead to constraint violations and, consequently, unsuccessful flexibility service delivery. This approach improves risk characterization and ensures the model effectively handles worst-case scenarios. An analytical solution is provided that is less computationally demanding compared to commonly used sample-based methods. Additionally, the study explores how varying failure thresholds set by Energinet impact the ancillary service market participation of the EV aggregator, offering practical insights for both aggregators and the TSO on risk assessment and supply of ancillary services.
@@ II Preliminaries
Figure 1: Illustrative examples of flexibility estimation: Figure (a) illustrates how downwards and upwards flexibility for the aggregated case are calculated. Downwards flexibility ($r^{\downarrow}$) is determined as the difference between the lowest point of the maximum consumption of the EVs and the highest actual consumption across the hour, while upwards flexibility ($r^{\uparrow}$) is the minimum of the actual consumption values during the hour. The figures in (b) show how energy flexibility ($r^{\rm{\rm{E_{20}}}}$) is calculated for a single EV and for the aggregated case. The single EV case represents a complete charging session, while the aggregated case captures a time interval in the day with large fluctuations in available flexibility, typically during the morning when many charging sessions end. Energy flexibility is calculated based on the current state of charge (SoC) and the amount of power that can be applied to the EV over the next 20 minutes without exceeding the battery’s capacity. For instance, at hour 1 in the single EV case, the SoC is 30 kWh, and the battery capacity is 75 kWh, meaning that, theoretically, the charge box (CB) could apply a maximum of 45 kWh in the next 20 minutes to fully charge the EV. However, due to the CB’s physical constraint of 12 kW, the actual available energy flexibility is much lower. Similarly, in the aggregated case, the theoretical and realistic energy flexibilities are on vastly different scales, reflecting the system’s capacity and constraints.
@@ II-B The P90 requirement
To maintain grid stability, the provision of ancillary services must be highly reliable. However, imposing a 100% reliability requirement for ancillary service bids would limit the participation of many non-conventional technologies, such as wind power or aggregated flexible loads like EVs. To facilitate the participation of stochastic flexible resources in the FCR markets, including FCR-D, which is the focus of this paper, Energinet has introduced the P90 requirement, as outlined in [2] and [5].
The P90 requirement ensures that flexibility providers, such as EV aggregators, can participate in the FCR markets, provided that their reserved capacity bid is fully realized at least 90% of the time. To be eligible to bid into the markets, an aggregator must complete Energinet’s prequalification process [2] and conduct a prognosis analysis of its ability to meet future demand, which is verified by Energinet. Once qualified by the TSO, a flexibility provider can participate in the relevant ancillary service markets. From that point, Energinet continuously performs ex-post checks to verify whether a reserved capacity bid by the participant is truly available, regardless of whether the reserve is activated. This serves as an additional security measure for supply. For example, an upwards reserve bid of 10 kW is considered unavailable if the actual consumption realization of an EV aggregator is only 8 kW. Any partial or full unavailability of a bid counts as a reserve shortfall. Energinet conducts these checks at appropriate time intervals, such as every quarter. If a provider violates the P90 requirement, it will be excluded from further market participation and must re-apply for qualification.
This requirement can be expressed mathematically using chance constraints of the form $\mathbb{P}(\cdot)\geq 1-\varepsilon$, where $\varepsilon=0.1$ represents the maximum 10% allowance for failure in successfully delivering the service.
@@ II-C The LER requirement
In addition to the P90 requirement, Energinet has introduced an additional requirement for resources with limited energy reservoirs (LER), such as batteries and EV aggregators. The primary objective of the LER requirement is to ensure that resources with limited energy capacity can maintain energy availability despite their constraints. This includes adhering to energy management protocols and ensuring the ability to sustain continuous activation for a specified minimum duration when needed. The definition of this requirement is outlined in [2] and [5].
The specifics of the LER requirement differ between the FCR-N and FCR-D services. This paper focuses on FCR-D services, where the requirement states that an LER unit bidding in one direction of the FCR-D market (either up or down) must also ensure that at least 20% of the reserve bid is available in the opposite direction. For EV aggregators, this LER requirement applies only in the downwards direction, since activation in this direction (i.e., increasing power consumption) is constrained by the battery energy storage capacity of EVs. In contrast, activation in the upwards direction (i.e., reducing power consumption) is not constrained by battery capacity and is treated similarly to other non-LER flexible loads. Additionally, the requirement mandates that the service must be available for at least 20 continuous minutes once activated.
These conditions impose the following constraints on the upwards and downwards reserve capacity bids of the EV aggregator (in kW), denoted as $b^{\uparrow}$ and $b^{\downarrow}$, which are submitted to the FCR-D market:
@@
where $\{R^{\uparrow},R^{\downarrow},R^{\rm{E_{20}}}\}$ are random variables representing the available upwards, downwards, and energy flexibility, respectively. Downwards flexibility, $R^{\downarrow}$, refers to the potential for increasing consumption of the grid-connected EV fleet, while upwards flexibility, $R^{\uparrow}$, refers to the potential for reducing consumption. In addition, energy flexibility, $R^{\rm{E_{20}}}$, is determined based on the current state of charge of the entire EV fleet and the maximum power that can be applied to the fleet over the next 20 minutes without exceeding battery capacity.
The realizations of these random variables are denoted as $\{r^{\uparrow},r^{\downarrow},r^{\rm{E_{20}}}\}$. A detailed explanation of how these realizations are defined can be found in Appendix A. For a quicker overview, Fig. 1 provides graphical representations of how the flexibilities are estimated. Specifically, the realizations $\{r^{\uparrow},r^{\downarrow},r^{\rm{E_{20}}}\}$ are defined as the extreme values of flexibility within the hour, which in this study corresponds to the minimum available flexibility during the hour. This approach differs from [5], where similar sets are defined using minute-level realizations. By considering the minimum value over the entire hour, the computational burden is reduced, as the number of data points decreases by a factor of 60.
II-D Case study
Our case study utilizes real-world data comprising charging profiles from 1400 residential EV charging boxes in Denmark. The measurements were recorded from March 24, 2022, to March 21, 2023, with an average time step of 2.84 minutes. To achieve a uniform resolution, the data has been interpolated to a one-minute granularity. It is assumed that each charging box is associated with a single EV. The historical EV consumption levels serve as the baseline for estimating each flexibility. Further details are provided in Appendix A.
@@ III Chance-constrained bidding model
To incorporate the requirements imposed by Energinet into the bidding strategy, we employ chance-constrained optimization. This approach manages uncertainty by ensuring that certain constraints are satisfied with a specified probability rather than absolute certainty. For EV aggregators participating in ancillary service markets, chance constraints enable bids that provide a balance between reliability and resource utilization. Specifically, Energinet’s grid code mandates that aggregators commit to a service provision that is fully available at least 90% of the time, allowing for a 10% margin of non-compliance due to uncertainties in EV availability.
In line with the P90 and LER requirements, we adopt a joint chance-constrained optimization model similar to the one proposed in [5]:
@@
where the objective function (2a) maximizes the combined reserve capacity bids into the FCR-D up and down markets. Additionally, (2b) defines the joint chance constraint that incorporates both the LER and P90 requirements. Specifically, the chance constraint $\mathbb{P}(\cdot)\geq 1-\varepsilon$ ensures that the probability of jointly satisfying the constraints within the corresponding hour is at least $1-\varepsilon$. In accordance with the P90 requirement, $\varepsilon=0.1$. Note that the bids are submitted on an hourly basis, and the proposed model accounts only for the reservation of flexibility capacity. As a result, the rebound effect is not considered, meaning there is no inter-hour dependency in the bidding scheme. Future work could address this by incorporating activation data.
@@
Chance-constrained programs are generally computationally intractable. However, if the underlying probability distribution satisfies certain properties, analytical reformulations of the chance constraints can be derived. For example, analytical methods like the Bonferroni approximation [15] transform probabilistic constraints into a series of more tractable constraints, simplifying the problem at the potential cost of precision. Alternatively, sample-based methods represent uncertainty through a set of likely scenarios, allowing for optimization that accounts for a range of possible outcomes. These approaches enable EV aggregators to optimize their bidding strategies while managing the stochastic nature of their resources within the operational constraints of ancillary service markets.
@@
In Fig. 2, scatter plots of the in-sample and out-of-sample data used in our case study are shown. In-sample data refers to the set used in the optimization problem to determine optimal bids, while out-of-sample data is used ex-post for bid evaluation. Extreme value theory [16] focuses on the behavior of extreme values (maxima or minima) within a dataset, particularly as the sample size increases. We are particularly interested in the extreme values, which in our case correspond to those below the dotted line in Fig. 2, representing the empirical 10th percentile of the in-sample data. To avoid bias, we solve the model in multiple runs, each with different sets of in-sample and out-of-sample data, while keeping the number of samples in each set unchanged.
@@
To derive an analytical solution to (2), we aim to fit a distribution to the tail data (samples below the empirical 10th percentile) that facilitates an analytical reformulation of the chance constraint. Among the various options for modeling tail data, we tested both the Weibull and Generalized Pareto distributions. We observed better out-of-sample performance with the former, and thus continue our analysis using the Weibull distribution.
@@
It is worth noting that in the literature on extreme value theory, the generalized extreme value distribution is more commonly used, as it emerges as the limiting distribution of the maximum of $n$ independent and identically distributed random variables when $n$ grows large. However, the convergence to this limiting distribution may be poor. Additionally, the generalized extreme value distribution introduces extra parameters that require tuning, which may also demand more samples.
@@ IV-A Fitting a Weibull distribution to the tail data
Aiming to obtain an analytical solution, we fit a Weibull distribution to the tails of our data. The rationale for choosing this distribution lies in its flexible shape parameters, allowing to capture a wide range of tail behaviors. This flexibility is particularly advantageous, as this distribution can effectively model both skewed and heavy-tailed data. Its ability to adapt to both light and heavy tails enables us to accurately capture the realistic variability and potential extremes in the data. The Weibull distribution has extensive applications in fields such as reliability engineering, where it models failure times and lifetimes, and in risk management for assessing minimum extreme events.
@@
which we maximize over a range of $\gamma$-values. Looking back at Fig. 3, the Weibull distribution is fitted to the lower tail of the sampled data, with the shape parameter $\hat{\gamma}=0.9265$. As it is needed in later analysis, we define the cumulative distribution function (CDF) of the Weibull distribution as: